Smart TVs have seen success in the United States market. During the 2nd quarter of
2015 45% of TVs sold in the United States were Smart TVs. That’s an increase of 11% from 2014.
Choose three households and find the probability that
(a) None of the 3 households had a Smart TV
(b) All 3 households had a Smart TV
(c) At least 1 of the 3 households had a Smart TV
1
Expert's answer
2020-10-23T10:57:43-0400

p=0.45

q=1-p=0.55

(a) None of the 3 households had a Smart TV

P=0.553=0.166

(b) All 3 households had a Smart TV

P=0.453=0.091

c) At least 1 of the 3 households had a Smart TV

P=1-P( None of the 3 households had a Smart TV)=1-0.166=0.834